Examine the data in the appendix to Chapter 19. Locate the figures for nominal and real GDP for 2006 and 2007. Calculate the GDP deflator. What were the rates of growth of nominal GDP and real GDP for 2007? What was the rate of inflation (as measured by the GDP deflator) for 2007? Compare the rate of inflation using the GDP deflator with that using the CPI.



1
Expert's answer
2021-09-04T18:22:16-0400

"GDP \space deflator=\frac{nominal GDP}{real \space GDP}\times 100"


"2007\\=\frac{12000}{10000}\times100=120\%\\2006\\=\frac{10000}{8000}\times100=125\%"


rate of growth of

i)nominal GDP in 2007

"=\frac{nominal\space GDP\space in\space 2007-nominal\space GDP\space in\space 2006}{nominal \space GDP\space in\space 2006}\times 100\\=\frac{12000-10000}{10000}\times 100\\=20\%"


ii)real GDP in 2007

"=\frac{real\space GDP\space in\space 2007-real\space GDP\space in\space 2006}{real \space GDP\space in\space 2006}\times 100\\=\frac{10000-8000}{8000}\times 100\\=25\%"


rate of inflation in 2007

"\frac{deflator\space in\space 2006-deflator \space in \space 2006}{deflator\space in\space 2006}\times 100\\=\frac{120-125}{125}\times 100\\=-4\%"

in GDP deflator, inflation measures level of prices of purchase by government, consumers and business while in CPI the inflation measures level of prices of purchase by consumers only.